OpenStudy (anonymous):

solve by elimination 3x + 4y=5 6x +7y =8

OpenStudy (anonymous):

Multiply the first equation by -2. Then add the 2 equations together and the x's will go away.

OpenStudy (anonymous):

-2 * (3x + 4y=5)<---------multiply the first equation by -2 -6x - 8y = -10 <------------now add this to the other equation 6x + 7y = 8 <------------the other equation --------------- 15 y = -2 Solve for y. Then substitute that into one of the equations and solve for x.